# GS-SDF: LiDAR-Augmented Gaussian Splatting and Neural SDF for Geometrically Consistent Rendering and Reconstruction

### v1

@@ Abstract
Digital twins are fundamental to the development of autonomous driving and embodied artificial intelligence. However, achieving high-granularity surface reconstruction and high-fidelity rendering remains a challenge. Gaussian splatting offers efficient photorealistic rendering but struggles with geometric inconsistencies due to fragmented primitives and sparse observational data in robotics applications. Existing regularization methods, which rely on render-derived constraints, often fail in complex environments. Moreover, effectively integrating sparse LiDAR data with Gaussian splatting remains challenging. We propose a unified LiDAR-visual system that synergizes Gaussian splatting with a neural signed distance field. The accurate LiDAR point clouds enable a trained neural signed distance field to offer a manifold geometry field, This motivates us to offer an SDF-based Gaussian initialization for physically grounded primitive placement and a comprehensive geometric regularization for geometrically consistent rendering and reconstruction. Experiments demonstrate superior reconstruction accuracy and rendering quality across diverse trajectories. To benefit the community, the codes will be released at https://github.com/hku-mars/GS-SDF.
@@ I Introduction
3D surface reconstructions and photorealistic renderings are essential for a wide range of applications, like augmented reality[1], and embodied artificial intelligence[2]. The increasing accessibility of cameras and cost-effective LiDAR sensors on robots has enabled the collection of rich multimodal data. LiDAR-visual SLAM technologies[3, 4, 5] are widely applied in 3D reconstruction tasks. However, these methods typically produce only discrete colorized raw point clouds, which lacks complete surface structure and rendering ability for view-dependent novel view synthesis. In practical digital-twin applications [6], high-quality watertight surface reconstructions and photorealistic rendering are critical.
@@
Neural Radiance Fields (NeRFs) [7] have gained attention for their ability to generate high-fidelity renderings. However, their computational intensity and inefficiency in real-time applications limit their practicality. 3D Gaussian Splatting (3DGS) [8] has emerged as a compelling alternative to NeRFs for efficient, high-quality rendering. Unlike NeRF requiring laborious volume rendering, 3DGS employs explicit ellipsoidal Gaussian primitives for efficient rasterization. However, the fragmented structure of these primitives compromises structural continuity and often misaligns with underlying geometry. This misalignment hinders geometrically consistent rendering, particularly in robotics applications where free-view trajectories yield insufficient observations. Furthermore, under image-only supervision with imperfect camera poses, 3DGS frequently overfits to photometric cues, producing floating artifacts. It motivates us to introduce geometric regularization into the 3DGS framework.
@@
Geometric inconsistencies in 3DGS manifest as rendering distortions. Current regularization strategies focus on rendering-depth-derived constraints, such as rendering normal consistency from 2D Gaussian splatting (2DGS) [9] or multi-view alignment from PGSR[10]. They focus on object-centric scenes with abundant multi-view regularization and limited generalization to complex environments. Depth cameras and LiDARs providing direct structural priors inspire us to incorporate these sensors into the 3DGS framework for geometric regularization. And depth cameras with limited precision are confined for indoor applications[11]. While LiDAR providing sparse point clouds remains challenging to integrate with NeRF[12] or 3DGS[13] frameworks designed for dense, image-scale geometry.
@@
The combination of LiDAR and neural signed distance fields (NSDF) has shown promise in high-granularity surface reconstructions[14]. The NSDF, providing a manifold geometry field, motivates us to explore a more comprehensive geometric regularization for Gaussian splats and the importance of the Gaussian initialization to the results. While, surface reconstruction from Gaussian splats is complicated by their discontinuous nature. The truncated Signed Distance Field (TSDF) fusion [15] is commonly applied for 3DGS-based surface reconstructions through depth images rendered from Gaussian splatting, which demonstrates unsatisfactory inaccuracies in surface reconstruction for generalized scenes.
@@
To address above limitations, we propose the combination of 2D Gaussian splatting and neural signed distance field for LiDAR-visual systems. We aim to reconstruct both the appearance and surface of generalized scenes using posed images and low-cost LiDAR data from any casual trajectories. We emphasize the importance role of the Gaussian initialization in returning a good structure of the Gaussian splats. We explore the combination of the NSDF and Gaussian splats to leverage the structural and photometric hints for high-granularity surface reconstruction and geometrically consistent rendering, as shown in Fig. 1. Our contributions are as follows:
@@
    A physically grounded Gaussian initialization based on the neural signed distance field to boost training convergency and reduce floating artifacts.
@@ II-A Geometrically Consistent Novel View Synthesis
NeRFs [7] and 3DGS [8] have been widely employed for novel view synthesis, owing to their remarkable photorealistic rendering capabilities. However, neither of them can guarantee geometric consistency, frequently leading to the emergence of floating artifacts [16]. To tackle this problem, geometric regularizations and structural priors derived from depth sensors have been incorporated into both NeRF and 3DGS. 2DGS [9] projects Gaussian ellipsoids into surfels and enforces normal-depth alignment geometric regularization. PGSR [10] provides further guarantees for multi-view rendering consistency regularization. RGBD cameras that provide aligned color and depth information are well practiced in both NeRFs [17, 18] and 3DGS [11] to align rendered depth images with sensor-acquired ones. However the low measurement accuracy of depth cameras limited their application in indoor scene settings. LiDAR sensors provide accurate depth information, but their sparse nature complicates integration with NeRFs [19] or 3DGS [20] frameworks designed for dense, image-scale geometry. M2Mapping [12] enforces geometric consistency to NeRF through the combination of NSDF and ray marching, albeit at a significant computational expense. LIV-GaussMap [20] stabilizes splat orientations using planar priors derived from LiDAR-inertial odometry. And LI-GS [13] utilizes Gaussian Mixture Models for constraining the Gaussian splats’ distribution.
@@ II-B Surface Reconstruction
Surface reconstructions are predominated using point clouds with direct structural information. The TSDF fusion [15] is well practiced for surface reconstructions[21]. While the quest for continuous surface reconstruction motivates the exploration of implicit representations, like Poisson functions [22] or NSDFs [23, 24]. In terms of the exploration of Gaussian splats for surface reconstruction, its discontinuous nature complicates surface reconstruction. Recent approaches circumvent this limitation through intermediate volumetric representations. SuGaR [25] computes an approximate density field based on the opacities of neighboring Gaussians, and GOF [26] defines per-point opacity as the minimal opacity across training views. However, both density and opacity fields suffer from computational inefficiency and ambiguous surface level sets for surface reconstruction. In contrast, the SDF facilitates precise surface extraction through defined zero-level sets. The TSDF fusion driven by the rendered depth images is favored by 3DGS-based surface reconstructions[9, 10], but yields insufficient accuracy. Image-driven NSDFs combined with 3DGS [27, 28] are still under exploration for achieving considerable accuracy in surface reconstructions.
@@ III Methodology
Given posed images and point clouds from LiDAR-inertial-visual odometry (e.g., FAST-LIVO2 [5]), we propose a framework to jointly reconstruct high-granularity scene geometry and photorealistic appearance. As shown in Fig. 2, our pipeline consists of three stages: A neural signed distance field (NSDF) is first trained using point clouds to establish a manifold geometry field (Sec. III-B). Gaussian splats are initialized from the NSDF (Sec. III-C) and a thorough SDF-aided shape regularization for both Gaussian splats and the NSDF (Sec. III-D) achieve a geometric-consistent rendering and reconstruction.
@@
$\mathcal{L}_{c}=0.8L_{1}({\boldsymbol{C}},\bar{\boldsymbol{C}})+0.2L_{\mathrm{DSSIM}}({\boldsymbol{C}},\bar{\boldsymbol{C}}),$ (2)
@@ III-B Neural Signed Distance Field
where $o_{i}=\Phi(-{s}_{i},{\beta}_{i})$ is the predicted occupancy value. The Eikonal regularization [30] is empolyed to ensure the gradient of the SDF maintains unit magnitude:
@@ III-C Initialization of Gaussian Splats
In this section, we address the crucial role of the Gaussian initialization plays a in optimizations and propose a careful Gaussian initialization with the help of a coarse neural signed distance field trained using input point cloud. We provide a thorough initialization of Gaussians boosts the convergency efficiency and suppresses floating artifacts.
@@ III-C3 Color Initialization
Next, we maintain fixed Gaussians’ structural attributes (position, rotation, scale, and opacity) while training the Gaussians on all training images in a single round for color initialization. The color initialization helps to solid the Gaussians’ structural from being deviated in early training stages.
@@ III-D Geometric Regularization for Rendering
The rasterization-based rendering of Gausisan splatting can lead to occlusion artifacts where foreground elements inappropriately obscure background content. Geometric regularization are designed to hold the rendering consistency of different views to avoid such artifacts. Considering that the LiDAR structure priors are not always available everywhere, we incorporate both the rendering-based regularization in 2D space and the SDF-aided shape regularization in 3D space to guarantee a geometrically consistent rendering.
@@ III-D1 Rendering Consistency Regularization
Following with the rendering regularization in 2DGS[9], we obtain the rendered normal images from splats’ normals: $\boldsymbol{N}=\sum_{i=1}\boldsymbol{n}_{i}w_{i}$. And the finite differential normal images from the rendered depth images: $\hat{\boldsymbol{N}}(x,y)=\frac{\nabla_{x}\boldsymbol{p}_{d}\times\nabla_{y}\boldsymbol{p}_{d}}{\|\nabla_{x}\boldsymbol{p}_{d}\times\nabla_{y}\boldsymbol{p}_{d}\|}$, where $\boldsymbol{p}_{d}$ is the point derived from the rendered depth images at piexl $(x,y)$. The rendering consistency regularization aligns the rendered normal with the finite differential normal as follows:
@@ III-D2 SDF-aided Shape Regularization
With the NSDF providing a manifold geometry field, we can regularize the Gaussian splats’ shape to align with the physical surface. Consider that the splats represent a local plane of the surface, simply pulling the splat center $\boldsymbol{p}$ to the zero-level sets is insufficient[28], as shown in Fig. 3(a). We propose the shape regularization to enables comprehensive structural regularization across the entire splat surface. For each splat, we sample a point $\boldsymbol{u}_{s}=[u_{s},v_{s}]^{T}\sim\mathcal{N}(0,1)$ on its disk surface:
@@
where $W_{i}=\sum_{k}w_{k}$ represents the accumulated weight of the $i$-th splat from Eq. 1. Considering that the SDF gradient maintains unit magnitude $\|\frac{\partial f(\boldsymbol{p}_{s_{i}})}{\partial\boldsymbol{p}_{s_{i}}}\|=1$, we adopt the $L_{2}$ loss function to enable adaptive refinement with proper gradients: $\frac{\partial\mathcal{L}_{s}}{\partial\boldsymbol{p}_{s_{i}}}=W_{i}\mathcal{G}(\boldsymbol{u}_{s_{i}})f(\boldsymbol{p}_{s_{i}})\frac{\partial f(\boldsymbol{p}_{s_{i}})}{\partial\boldsymbol{p}_{s_{i}}}$.
@@
Figure 3: Illustration of different geometric regularization. The blue splats represent the initial splat, and the orange splats represent the optimized splat. The dashed lines represent the optimization directions.
@@ III-E Optimization
The overall optimization function is a combination of the supervisions from point clouds, images, and geometric regularizations:
@@ IV-A Implementation Details
We represent our neural signed distance field following a similar architecture to InstantNGP [29], utilizing a combination of multi-resolution hash encoding and a tiny MLP decoder. Given any position $x$, the hash encoding concatenates each resolution’s interpolation features to form a feature vector of size $32$. A geometry MLP with 64-width and 3 hidden layers decodes the encoding feature to obtain the SDF value and scale. We sample 4096 rays for NSDF training with 4 surface samples and 4 free space samples on each ray. We first take 10000 iterations for NSDF training and leverage the trained NSDF for our initialization as described in Sec. III-C. Then, we follow the setting of 2DGS[9] to take 30000 iterations for the training. The implementation is based on LibTorch and CUDA, and all experiments are conducted on a platform equipped with an Intel i7-13700K CPU and an NVIDIA RTX 4090 GPU.
@@
Figure 4: Reconstructed mesh on the Replica dataset’s Office-2 and colors indicate the direction of the surface normal. Our method can capture more precise geometric details in slim objects.
@@
Figure 5: Extrapolation rendering results on the Replica dataset’s room-0. Our method can capture more precise geometric details in slim objects and retain both the structure and texture in extrapolation rendering.
@@ IV-B2 Metrics
To evaluate the quality of the geometry, we recover the NSDF to a triangular mesh using marching cubes [31] and calculate the Chamfer Distance (C-L1, cm) and F-Score ($<2$ cm, %) with the ground truth. For rendering quality evaluation, we use the Structural Similarity Index (SSIM) and Peak Signal-to-Noise Ratio (PSNR) to compare the rendered image with the ground truth image.
@@ IV-C1 Replica Datasets
The Replica datasets [6] provides room-scale indoor simulation RGBD sensor data. We follow the M2Mapping[12] to emphasize the issues of extrapolation rendering consistency with extrapolation evaluation datasets generated by uniformly sampling positions and orientations in each scene from Replica. The quantitative resutls, comparing surface reconstruction, interpolation, and extrapolation rendering, are shown in Tab. I, our method shows the competitive surface reconstruction results with the state of the art, and the best performance in terms of interpolation rendering but fails to compete with the extrapolation rendering results of the NeRF-based M2Mapping. We validate our method is capable of capturing more precise geometric details in slim objects, as shown in Fig. 4. In the extrapolation views shown in Fig. 5, our method can retain more high-frequency textures (sharper spots) while NeRF-based methods show more natural transition (smoother lighting).

### v2

@@ Abstract
Digital twins are fundamental to the development of autonomous driving and embodied artificial intelligence. However, achieving high-granularity surface reconstruction and high-fidelity rendering remains a challenge. Gaussian splatting offers efficient photorealistic rendering but struggles with geometric inconsistencies due to fragmented primitives and sparse observational data in robotics applications. Existing regularization methods, which rely on render-derived constraints, often fail in complex environments. Moreover, effectively integrating sparse LiDAR data with Gaussian splatting remains challenging. We propose a unified LiDAR-visual system that synergizes Gaussian splatting with a neural signed distance field. The accurate LiDAR point clouds enable a trained neural signed distance field to offer a manifold geometry field. This motivates us to offer an SDF-based Gaussian initialization for physically grounded primitive placement and a comprehensive geometric regularization for geometrically consistent rendering and reconstruction. Experiments demonstrate superior reconstruction accuracy and rendering quality across diverse trajectories. To benefit the community, the codes are released at https://github.com/hku-mars/GS-SDF.
@@ I Introduction
3D surface reconstructions and photorealistic renderings are essential for a wide range of applications, like augmented reality[1], and embodied artificial intelligence[2]. The increasing accessibility of cameras and cost-effective LiDAR sensors on robots has enabled the collection of rich multimodal data. LiDAR-visual SLAM technologies[3, 4, 5] are widely applied in 3D reconstruction tasks. However, these methods typically produce only discrete colorized raw point clouds, which lack complete surface structure and rendering ability for view-dependent novel view synthesis. In practical digital-twin applications [6], high-quality watertight surface reconstructions and photorealistic rendering are critical.
@@
Neural Radiance Fields (NeRFs) [7] have gained attention for their ability to generate high-fidelity renderings. However, their computational intensity and inefficiency in real-time applications limit their practicality. 3D Gaussian Splatting (3DGS) [8] has emerged as a compelling alternative to NeRFs for efficient, high-quality rendering. Unlike NeRF, which requires laborious volume rendering, 3DGS employs explicit ellipsoidal Gaussian primitives for efficient rasterization. However, the fragmented structure of these primitives compromises structural continuity and often misaligns with underlying geometry. This misalignment hinders geometrically consistent rendering, particularly in robotics applications where free-view trajectories yield insufficient observations. Furthermore, under image-only supervision with imperfect camera poses, 3DGS frequently overfits to photometric cues, producing floating artifacts. It motivates us to introduce geometric regularization into the 3DGS framework.
@@
Geometric inconsistencies in 3DGS manifest as rendering distortions. Current regularization strategies focus on rendering-depth-derived constraints, such as rendering normal consistency from 2D Gaussian splatting (2DGS) [9] or multi-view alignment from PGSR[10]. They focus on object-centric scenes with abundant multi-view regularization and limited generalization to complex environments. Depth cameras and LiDARs providing direct structural priors inspire us to incorporate these sensors into the 3DGS framework for geometric regularization. And depth cameras with limited precision are confined to indoor applications[11]. While LiDAR provides sparse point clouds, it remains challenging to integrate with NeRF[12] or 3DGS[13] frameworks designed for dense, image-scale geometry.
@@
The combination of LiDAR and neural signed distance fields (NSDF) has shown promise in high-granularity surface reconstructions[14]. The NSDF, providing a manifold geometry field, motivates us to explore a more comprehensive geometric regularization for Gaussian splats and the importance of the Gaussian initialization to the results. Surface reconstruction from Gaussian splats is complicated by their discontinuous nature. The truncated Signed Distance Field (TSDF) fusion [15] is commonly applied for 3DGS-based surface reconstructions through depth images rendered from Gaussian splatting, which demonstrates unsatisfactory inaccuracies in surface reconstruction for generalized scenes.
@@
To address the above limitations, we propose the combination of 2D Gaussian splatting and neural signed distance field for LiDAR-visual systems. We aim to reconstruct both the appearance and surface of generalized scenes using posed images and low-cost LiDAR data from any casual trajectories. We emphasize the important role of the Gaussian initialization in returning a good structure of the Gaussian splats. We explore the combination of the NSDF and Gaussian splats to leverage the structural and photometric hints for high-granularity surface reconstruction and geometrically consistent rendering, as shown in Fig. 1. Our contributions are as follows:
@@
    A physically grounded Gaussian initialization based on the neural signed distance field to boost training convergence and reduce floating artifacts.
@@ II-A Geometrically Consistent Novel View Synthesis
NeRFs [7] and 3DGS [8] have been widely employed for novel view synthesis, owing to their remarkable photorealistic rendering capabilities. However, neither of them can guarantee geometric consistency, frequently leading to the emergence of floating artifacts [16]. To tackle this problem, geometric regularizations and structural priors derived from depth sensors have been incorporated into both NeRF and 3DGS. 2DGS [9] projects Gaussian ellipsoids into surfels and enforces normal-depth alignment geometric regularization. PGSR [10] provides further guarantees for multi-view rendering consistency regularization. RGBD cameras that provide aligned color and depth information are well practiced in both NeRFs [17, 18] and 3DGS [11] to align rendered depth images with sensor-acquired ones. However, the low measurement accuracy of depth cameras limited their application in indoor scene settings. LiDAR sensors provide accurate depth information, but their sparse nature complicates integration with NeRFs [19] or 3DGS [20] frameworks designed for dense, image-scale geometry. M2Mapping [12] enforces geometric consistency to NeRF through the combination of NSDF and ray marching, albeit at a significant computational expense. LIV-GaussMap [20] stabilizes splat orientations using planar priors derived from LiDAR-inertial odometry. And LI-GS [13] utilizes Gaussian Mixture Models for constraining the Gaussian splats’ distribution.
@@ II-B Surface Reconstruction
Surface reconstructions predominantly use point clouds with direct structural information. The TSDF fusion [15] is well practiced for surface reconstructions[21]. While the quest for continuous surface reconstruction motivates the exploration of implicit representations, like Poisson functions [22] or NSDFs [23, 24]. In terms of the exploration of Gaussian splats for surface reconstruction, their discontinuous nature complicates surface reconstruction. Recent approaches circumvent this limitation through intermediate volumetric representations. SuGaR [25] computes an approximate density field based on the opacities of neighboring Gaussians, and GOF [26] defines per-point opacity as the minimal opacity across training views. However, both density and opacity fields suffer from computational inefficiency and ambiguous surface-level sets for surface reconstruction. In contrast, the SDF facilitates precise surface extraction through defined zero-level sets. The TSDF fusion driven by the rendered depth images is favored by 3DGS-based surface reconstructions[9, 10], but yields insufficient accuracy. Image-driven NSDFs combined with 3DGS [27, 28] are still under exploration for achieving considerable accuracy in surface reconstructions.
@@ III Methodology
Given posed images and point clouds from LiDAR-inertial-visual odometry (e.g., FAST-LIVO2 [5]), we propose a framework to jointly reconstruct high-granularity scene geometry and photorealistic appearance. As shown in Fig. 2, our pipeline consists of three stages: A neural signed distance field (NSDF) is first trained using point clouds to establish a manifold geometry field (Sec. III-B). Gaussian splats are initialized from the NSDF (Sec. III-C), and a thorough SDF-aided shape regularization for both Gaussian splats and the NSDF (Sec. III-D) achieves a geometric-consistent rendering and reconstruction.
@@
[…]
@@ III-B Neural Signed Distance Field
where $o_{i}=\Phi(-{s}_{i},{\beta}_{i})$ is the predicted occupancy value. The Eikonal regularization [30] is employed to ensure the gradient of the SDF maintains unit magnitude:
@@ III-C Initialization of Gaussian Splats
In this section, we address the crucial role that Gaussian initialization plays in optimizations and propose a careful Gaussian initialization with the help of a coarse neural signed distance field trained using an input point cloud. We provide a thorough initialization of Gaussians, which boosts the convergence efficiency and suppresses floating artifacts.
@@ III-C3 Color Initialization
Next, we maintain fixed Gaussians’ structural attributes (position, rotation, scale, and opacity) while training the Gaussians on all training images in a single round for color initialization. The color initialization helps to solidify the Gaussians’ structure from being deviated in early training stages.
@@ III-D Geometric Regularization for Rendering
The rasterization-based rendering of Gaussian splatting can lead to occlusion artifacts where foreground elements inappropriately obscure background content. Geometric regularization is designed to hold the rendering consistency of different views to avoid such artifacts. Considering that the LiDAR structure priors are not always available everywhere, we incorporate both the rendering-based regularization in 2D space and the SDF-aided shape regularization in 3D space to guarantee a geometrically consistent rendering.
@@ III-D1 Rendering Consistency Regularization
[…]
@@ III-D2 SDF-aided Shape Regularization
With the NSDF providing a manifold geometry field, we can regularize the Gaussian splats’ shape to align with the physical surface. Consider that the splats represent a local plane of the surface, simply pulling the splat center $\boldsymbol{p}$ to the zero-level sets is insufficient[28], as shown in Fig. 3(a). We propose the shape regularization to enable comprehensive structural regularization across the entire splat surface. For each splat, we sample a point $\boldsymbol{u}_{s}=[u_{s},v_{s}]^{T}\sim\mathcal{N}(0,1)$ on its disk surface:
@@
where $W_{i}=\sum_{k}w_{k}$ represents the accumulated weight of the $i$-th splat from Eq. 1 in a image. Considering that the SDF gradient maintains unit magnitude $\|\frac{\partial f(\boldsymbol{p}_{s_{i}})}{\partial\boldsymbol{p}_{s_{i}}}\|=1$, we adopt the $L_{2}$ loss function to enable adaptive refinement with proper gradients: $\frac{\partial\mathcal{L}_{s}}{\partial\boldsymbol{p}_{s_{i}}}=W_{i}\mathcal{G}(\boldsymbol{u}_{s_{i}})f(\boldsymbol{p}_{s_{i}})\frac{\partial f(\boldsymbol{p}_{s_{i}})}{\partial\boldsymbol{p}_{s_{i}}}$.
@@
Figure 3: Illustration of different geometric regularizations. The blue splats represent the initial splat, and the orange splats represent the optimized splat. The dashed lines represent the optimization directions.
@@ III-E Optimization
The overall optimization function is a combination of the supervision from point clouds, images, and geometric regularizations:
@@ IV-A Implementation Details
We represent our neural signed distance field following a similar architecture to InstantNGP [29], utilizing a combination of multi-resolution hash encoding and a tiny MLP decoder. Given any position $x$, the hash encoding concatenates each resolution’s interpolation features to form a feature vector of size 323232. A geometry MLP with 64-width and 3 hidden layers decodes the encoding feature to obtain the SDF value and scale. We sample 4096 rays for NSDF training with 4 surface samples and 4 free space samples on each ray. We first take 10000 iterations for NSDF training and leverage the trained NSDF for our initialization as described in Sec. III-C. Then, we follow the setting of 2DGS[9] to take 30000 iterations for the training. The implementation is based on LibTorch and CUDA, and all experiments are conducted on a platform equipped with an Intel i7-13700K CPU and an NVIDIA RTX 4090 GPU.
@@
Figure 4: Reconstructed mesh on the Replica dataset’s Office-2 and colors indicate the direction of the surface normal. Our method and M2Mapping can capture precise geometric details in slim objects.
@@
Figure 5: Extrapolation rendering results on the Replica dataset’s room-0. Our method can preserve great structure and texture details, while the NeRF-based method, M2Mapping, retains better consistency on the left wall.
@@ IV-B2 Metrics
To evaluate the quality of the geometry, we recover the NSDF to a triangular mesh using marching cubes [31] and calculate the Chamfer Distance (C-L1, cm) and F-Score (<2<2< 2 cm, %) with the ground truth. For rendering quality evaluation, we use the Structural Similarity Index (SSIM) and Peak Signal-to-Noise Ratio (PSNR) to compare the rendered image with the ground truth image.
@@ IV-C1 Replica Datasets
The Replica datasets [6] provide room-scale indoor simulation RGBD sensor data. We follow the M2Mapping[12] to emphasize the issues of extrapolation rendering consistency with extrapolation evaluation datasets generated by uniformly sampling positions and orientations in each scene from Replica. The quantitative results, comparing surface reconstruction, interpolation, and extrapolation rendering, are shown in Tab. I. Our method shows competitive surface reconstruction results with the state of the art, and the best performance in terms of interpolation rendering, but fails to compete with the extrapolation rendering results of the NeRF-based M2Mapping. We validate that our method is capable of capturing more precise geometric details in slim objects, as shown in Fig. 4. In the extrapolation views shown in Fig. 5, our method can retain more high-frequency textures (sharper spots) while NeRF-based methods show more natural transitions (smoother lighting).